OpenStudy (anonymous):

Write an equation in point slope form for a line with a point at (4,-5) and a slope of 2.

jimthompson5910 (jim_thompson5910):

Similar Example: If you're given a slope of 3 and a point of (1,-2), then y - y1 = m(x - x1) ... Start with point-slope form y - y1 = 3(x - x1) ... Plug in the given slope m = 3 y - (-2) = 3(x - 1) ... Plug in the given point (1, -2) for x1 and y1 y + 2 = 3(x - 1) ... Simplify